The mean length of certain construction lumber is supposed to be 8.5 feet. A random

sample of 21 pieces of such lumbers gives a sample mean of 8.3 feet and a sample

standard deviation of 1.2 feet. A builder claims that the mean of the lumber is different

from 8.5 feet. Does the datasupport the builder's claim at a= 0.05?


Step:

1.State the null and alternative hypothesis concerning the population mean, "\mu" and the type of test to be used.

2.Specify the level of significance "\alpha"

3.State the decision rule

4.Collect the data and perform calculations.

5. Make a statistical decision.

6.State the conclusion.


1
Expert's answer
2022-05-08T14:05:25-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=8.5"

"H_a:\mu\not=8.5"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


2. Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=20" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.085963."


3. The rejection region for this two-tailed test is "R = \{t: |t| >2.085963\}."


4. The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{8.3-8.5}{1.2/\sqrt{21}}\approx-0.76376"


5. Since it is observed that "|t| =0.76376< 2.085963=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=20" degrees of freedom, "t=-0.76376" is "p=0.453927," and since "p= 0.453927>0.05=\alpha," it is concluded that the null hypothesis is not rejected.


6.Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 8.5, at the "\alpha = 0.05" significance level.